Divaseventeen:

Living room furniture with a cash price of $2,900 at $100.00 down and $89.50 per month for 36 months.

ImBadKidTae:

what do you think

Divaseventeen:

I dont know

Divaseventeen:

3,something

jhonyy9:

so what is the question here ?

jhonyy9:

@XioGonz any idea pls ?

Divaseventeen:

Their three parts to this question

XioGonz:

Look at it this way: You are paying a total of 36* 89.5=3222.0 (for the 36 months) Your downpayment was 100 The total number of the payment made will be= 3222+100 =3,322. So you would pay 3222-2900 =$322

Divaseventeen:

Thankyou

XioGonz:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Divaseventeen Thankyou \(\color{#0cbb34}{\text{End of Quote}}\) No prob

Divaseventeen:

😁😁

jhonyy9:

good job @XioGonz

XioGonz:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jhonyy9 good job @XioGonz \(\color{#0cbb34}{\text{End of Quote}}\) Thank you

jhonyy9:

yw